what is the volume of a cube if the area of the base is 625 cm squared

Answer by ewatrrr(24785)

 
Hi,
what is the volume of a cube if the area of the base is 625 cm^2
A = s^2 = 625cm^2, s = 25cm
V = s^3 = (25cm)^3 = 15,625cm^3